# Quantum-Classical Hybrid Quantized Neural Network

### v1

@@ 6.3 Robust of Hybrid Approach: Impact of Coefficient Matrix Rounding
Figure 7 displays the QUBO matrix evolution using circular heatmaps, highlighting the impact of precision levels on iterative trajectories. Despite the differences revealed in these trajectories, our results confirm that all trajectories converge to the optimal solution, consistent with the theoretical convergence of the QCGD algorithm. Notably, even with minimal precision (e.g., 1 digit), matrix changes is small in later iterations. Those minimal matrix changes in the iterations across all precisions suggest that a lazy implementation of QCGD is possible, potentially reducing redundant quantum calls and enhancing efficiency on quantum hardware.
Figure 8: The figure comprises two subplots showing the smoothed absolute differences $|\mathrm{obj}_{k}-\mathrm{obj}_{10}|$ (left) and $|\mathrm{residual}_{k}-\mathrm{residual}_{10}|$ (right) relative to the 10-digit precision baseline across iterations for precision levels of 1 to 9 digits. The y-axis is on a log scale, with dashed lines representing each precision level.
@@
The smoothed absolute differences of the objective value and constraint residual relative to the 10-digit precision baseline were recorded, with results visualized in Figure 8. This setup allowed us to assess the impact of numerical precision on convergence trajectories. Both subplots of Figure 8 demonstrate that differences across all precision levels remain small, typically typically less than $10^{-2}$, indicating the hybrid algorithm’s robustness to precision variations. A downward trend is observed in difference of constraint residual, with differences decreasing sharply after initial fluctuations and stabilizing post-1500 iterations, confirming convergence to a consistent optimal solution regardless of precision, as supported by prior results. The rapid initial decline in the constraint residual difference, particularly for lower precisions (e.g., 1-3 digits), suggests early sensitivity to numerical accuracy, while the differences remain stably small in later iterations. The near-overlap of higher precision curves (e.g., 7-9 digits) highlights diminishing returns from increased precision.
7 Conclusion
@@
In conclusion, this paper introduces a novel QUBO model supporting arbitrary activation and loss functions through spline interpolation, enhancing quantum computing’s capability in neural network training. The theoretical upper bound on ising spins requirements and the demonstrated convergence of the hybrid algorithm, even with random output and limited coefficient matrix precision, underscore the robustness of our approach. The comparison with classical methods demonstrates the advantage of the Ising machine in solving quantized neural network training problem, while the hybrid algorithm exhibits consistent effectiveness and robustness. Moving forward, our model holds promise for applications requiring trainable activation functions, opening new avenues for enhancing neural network architectures with quantum computing technologies. This study marks a significant step towards realizing the full potential of quantum-enhanced machine learning algorithms in practical settings.
Appendix A Additional Related Work
In addition to the aforementioned work, there are also some other works related to quantum computing and neural networks. [49] introduces a novel training approach employing Adiabatic Quantum Computing (AQC), leveraging adiabatic evolution principles to address optimization problems. This universal AQC method proposed in [49] is designed for implementation on gate quantum computers. A hybrid approach is proposed in [50], which combines quantum and classical methods for accelerated training of BNN. The quantum portion utilizes the HHL algorithm to solve linear systems of equations for linear regression in a single layer BNN. In [51], the primary contribution lies in demonstrating the transfer of a trained artificial neural network to a quantum computing setting. The authors design a quadratic binary model for quantum annealing, aligning it with the behavior of a classical neural network by combining deep-learned parameters and layer structures.
Appendix B Error Bound of Piecewise Linear Approximation in FNN
@@
Appendix C Piecewise Linear Segment
Piecewise linear fitting approximates the function $f(\cdot)$ using linear segments. This method can effectively capture the behavior of moderately nonlinear functions while maintaining a manageable level of complexity in the optimization problem. According to [52], by using binary variables, minimizing a piecewise linear functions $S(\cdot)$ can be represented in linear form:
$\min_{\alpha_{i}}$ $\sum_{i=0}^{n}{\alpha_{i}\cdot S(M_{i})}$ (59)
s.t $\sum_{i=0}^{n}\alpha_{i}=1$
$\alpha_{0}\leq\beta_{1},\alpha_{n}\leq\beta_{n}$
$\alpha_{i}\leq\beta_{i+1}+\beta_{i}\;(1\leq i\leq n-1)$
$\sum_{i=1}^{n}{\beta_{i}}=1,\beta_{i}\in\{0,1\}$ (60)
where $\alpha_{i}\in[0,1]$ and slope of the piecewise linear function changes at $M_{i}(1\leq i\leq n)$.
Note that in this formulation, when $M_{k-1}\leq x\leq M_{k}$ for some $k$, then there exists some number $\alpha_{k-1}\in[0,1]$ such that
$x=\alpha_{k-1}M_{k-1}+(1-\alpha_{k})M_{k}.$ (61)
Since $S(\cdot)$ is linear in $[M_{k-1},M_{k}]$, thus
$S(x)=\alpha_{k-1}S(M_{k-1})+(1-\alpha_{k})S(M_{k})$ (62)
We remark that in the context of QUBO modeling, the introduction of quadratic terms in the objective function allows for simplification by eliminating the need to explicitly represent the variable $\alpha_{i}$ using binary encoding. This optimization is particularly advantageous as it significantly reduces the required number of bits. In addition, the linear form presented earlier in the formulation introduces a notable drawback: the absence of the decision variables as an independent entities. Instead, the decisions are indirectly represented through the coefficient $\alpha_{i}$, signifying its position within a specified range. Consequently, an additional relationship between the decision variables and $\alpha_{i}$ needs to be established.
To address these considerations, an alternative formulation is proposed below. This formulation not only leverages the inclusion of quadratic terms but also rectifies the issue of $\mathbf{x}$ indirect representation:
$\min\sum_{i=1}^{n+1}$ ${\Big{(}{\beta_{i-1}}\cdot\frac{h(\mathbf{x})-M_{i-1}}{M_{i}-M_{i-1}}+{\beta_{i}}\cdot\frac{M_{i}-h(\mathbf{x})}{M_{i}-M_{i-1}}\Big{)}\cdot f(M_{i-1})}$ (63)
@@ Appendix C Piecewise Linear Segment
For notational convenience, we set $\beta_{0}=\beta_{n+1}=0$. In the aforementioned objective function, we leverage the property that if $h(\mathbf{x})$ resides within the interval $[M_{i-1},M_{i}]$, we approximate the values of $f(\cdot)$ within this interval using line segments connecting points $(M_{i-1},f(M_{i-1}))$ and $(M_{i},f(M_{i}))$ in the two-dimensional coordinates. The value of the objective function at the location of $h(\mathbf{x})$ along this linear segment is determined by the corresponding linear interpolation. Mathematically, the linear function representing the approximation of $f(\cdot)$ within the interval $[M_{i-1},M_{i}]$ is expressed as:
$S(h(\mathbf{x}))$ $=f(M_{i-1})+\frac{f(M_{i})-f(M_{i-1})}{M_{i}-M_{i-1}}\cdot(h(\mathbf{x})-M_{i-1})$
$=\frac{M_{i}-h(\mathbf{x})}{M_{i}-M_{i-1}}\cdot f(M_{i-1})+\frac{h(\mathbf{x})-M_{i-1}}{M_{i}-M_{i-1}}\cdot f(M_{i})$ (66)
We introduce the weighting factor $\beta_{i}$ before the term in (66) to indicate whether $h(\mathbf{x})$ lies within the $i$-th interval $[M_{i-1},M_{i}]$. The overall objective function aggregates these weighted values over all intervals, summarizing the importance of each interval’s contribution to the optimization model.
Appendix D Proof of Convergence
@@
Appendix E Analysis of Total Error Due to Truncation
@@
$\Delta\text{Hamiltonian}$ $=|E[\sigma|J_{ij},h_{i}]-E[\sigma^{*‘}|J_{ij},h_{i}]|=O(2^{-d}\cdot n^{2}),$ (90)
ΔQUBO $\Delta\text{QUBO Value}$ $=|f(\mathbf{x}|q_{ij})-f(\mathbf{x}^{*’}|q_{ij})|=O(2^{-d}\cdot n^{2}).$ (91)
@@
Appendix F Additional Details of Experimental Results
@@
Appendix G Principles of STE and BinaryConnect

### v2

@@ Results
Figure 5(a) displays the QUBO matrix evolution using circular heatmaps, highlighting the impact of precision levels on iterative trajectories. Despite the differences revealed in these trajectories, our results confirm that all trajectories converge to the optimal solution, consistent with the theoretical convergence of the QCGD algorithm. Notably, even under minimal precision conditions (e.g., 1-digit rounding), matrix adjustments in later iterations remain marginal. These consistently small perturbations across all precision levels suggest that a lazy evaluation strategy for QCGD (Quantum-Constrained Gradient Descent) could be viable. Such an approach would minimize redundant quantum circuit executions, thereby improving computational efficiency on quantum hardware.
@@
The smoothed absolute differences of the objective value and constraint residual relative to the 10-digit precision baseline were recorded, with results visualized in Figure 5(b) and 5(c). This setup allowed us to assess the impact of numerical precision on convergence trajectories. Both subplots demonstrate that differences across all precision levels remain small, typically less than $10^{-2}$, indicating the hybrid algorithm’s robustness to precision variations. A downward trend is observed in the difference of constraint residual, with differences decreasing sharply after initial fluctuations and stabilizing post-1500 iterations, confirming convergence to a consistent optimal solution regardless of precision, as supported by prior results. The rapid initial decline in the constraint residual difference, particularly for lower precisions (e.g., 1-3 digits), suggests early sensitivity to numerical accuracy, while the differences remain stably small in later iterations. The near-overlap of higher precision curves (e.g., 7-9 digits) highlights diminishing returns from increased precision.
Methods
Copositive Programming Framework
The QCGD method is build on copositive programming (CP) [51] which is a mathematical optimization branch that has garnered significant attention for its applications in various optimization problems. In copositive programming, the objective is to optimize over the cone of copositive matrices, which are matrices for which the quadratic form yields nonnegative values over the nonnegative orthant. One key application of copositive programming is in transforming binary quadratic problems into completely positive problems. This transformation involves reformulating the standard quadratic problem as a CP program, enabling the optimization of a quadratic form over the Cartesian product of simplices.
$\min\mathbf{x}^{T}\mathbf{Q}\mathbf{x}$ (33)
$s.t.\mathbf{A}\mathbf{x}=\mathbf{b}.$ (34)
According to [45, 51], the quadratic objective and linear constraints of the quadratic binary optimization problem can be represented as linear functions of a copositive matrix, and the problems (33)-(34) can be converted to the completely positive program.
$\min\;$ $\mathbf{x}^{T}\mathbf{Q}_{0}\mathbf{x}$ (35)
$s.t.\;$ $\mathbf{A}\mathbf{x}=\mathbf{b}$ (36)
$\mathbf{x}^{T}\mathbf{Q}_{t}\mathbf{x}\leq c_{t},1\leq t\leq T$ (37)
In addition, we have the following improved linearization procedure. Consider the term $z^{l}_{x,j,k}=w^{l}_{jk}a^{l-1}_{x,k}$ in QCBO model, which can be expressed as $z^{l}_{x,j,k}=\sum_{\ell=1}^{\bar{\ell}}2^{\ell}\cdot w^{l}_{jk}\cdot\delta^{(\ell)}_{a^{l}_{x,k}}$. In our neural network training model, quadratic constraints arise from the $w^{l}_{jk}a^{l-1}_{x,k}$ terms. Note that the bilinear components involving $w^{l}_{jk}$ and $a^{l-1}_{x,k}$ share the common variable $w^{l}_{jk}$. To linearize these terms effectively, we can use the following constraints:
$z^{l}_{x,j,k}\leq(2^{\bar{\ell}+1}-2)\cdot w^{l}_{jk}.$ (40)
Note that here we treat $w^{\ell}_{jk}$ as binary variables; for integer-valued $w^{\ell}_{jk}$, the linearization process described above is still applicable. We only need to perform the same operation on the bits of the binary encoding of $w^{\ell}_{jk}$ and $a^{l-1}_{x,k}$.
Quantum-Conditional Gradient Descent and Lazy Implementation
  1. 1.
  2. 2.
  3. 3.
    Form the gradient matrix $\mathbf{G}_{t}$ using the coefficient matrix in the copositive programming formulation and constraint errors above. To ensure convergence, solve the subproblem $\min_{\mathbf{w}\in\mathbb{Z}_{2}^{p}}\mathbf{w}^{\top}\mathbf{G}_{t}\mathbf{w}$ $m_{t}$ times on a quantum annealer, selecting the best solution as:
  4. 4.
  5. 5.
  6. 6.
The practical implementation of the QCGD algorithm is challenged by the error rates associated with quantum computers, leading to inaccuracies in the computation results. This inaccuracy is particularly pronounced in the oracle, which plays a crucial role in the QCGD algorithm by providing direction in each iteration. The consideration of the inexact oracle in the context of QCGD is therefore crucial for several reasons. Firstly, due to the inherent limitations of quantum hardware, achieving perfect precision in quantum computations is challenging. The inexactness in the oracle reflects these limitations and underscores the need to develop algorithms that are robust to such imperfections. Secondly, the use of an inexact oracle also introduces a trade-off between computational accuracy and efficiency.
Definition 1 ($(\delta,\varepsilon)$-Inexact oracle [52, 53])
For problem
$\min_{\mathbf{D}}\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}\mathbf{D})$ (42)
A quantum computer without any imperfections corresponds to the scenario where $\delta=1$ and $\varepsilon=0$. The following lemma indicates that convergence can still be achieved even if the single quantum computation acts as an inexact oracle. This theoretical framework ensures that the overall algorithm remains robust despite potential inaccuracies in the quantum computation. Theoretical proof follows from that in [45], and we include it in the supplementary for completeness.
Proposition 1
[…]
To reduce unnecessary computation, we can adopt a lazy implementation strategy, where solving the QUBO problem is applied only when significant changes occur. This approach is motivated by the observation that, in later iterations, the magnitude of QUBO coefficient modifications often becomes negligible. We can determine whether solving the QUBO problem is necessary based on the spectral gap:
Then, it implies that the perturbations in the solution landscape are insignificant. Consequently, the algorithm can bypass solving the latest QUBO problem. By leveraging this stability, potentially through a threshold-based criterion to detect negligible changes, computational overhead can be reduced, particularly in resource-constrained environments.
Conclusion
@@
In conclusion, this paper introduces a QUBO-based framework that enables arbitrary activation and loss functions through spline interpolation, significantly expanding the applicability of quantum computing in neural network training. The theoretically derived upper bound on Ising spin requirements, along with the empirically validated convergence of our hybrid algorithm, even under randomized outputs and limited coefficient matrix precision, further validates the robustness of our approach. The comparison with classical methods highlights the superior performance of the Ising machine in quantized neural network training, while the hybrid algorithm demonstrates consistent efficacy and stability. Meanwhile, our model offers promising potential for applications demanding trainable activation functions, paving the way for quantum-enhanced neural architectures. This work represents a pivotal advancement toward harnessing the full potential of quantum-accelerated machine learning in real-world implementations.
Supplementary Material
Appendix A Error Bound of Piecewise Linear Approximation in FNN
@@
Appendix B Spline Quantumization Protocol
Formally we consider the following problem:
$\min_{\mathbf{x}\in\{0,1\}^{n}}f(h(\mathbf{x})),$ (57)
    $S(h(\mathbf{x}))=\sum_{i=1}^{n}S_{i}(h(\mathbf{x}))\cdot\beta_{i},$ (58)
B.1 Piecewise Constant Segment
$\min$ $\sum_{i=0}^{n}{{\beta_{i}}\cdot S(M_{i})}$ (59)
@@ B.1 Piecewise Constant Segment
Dealing with the inequality constraint.
We have now derived a QUBO model capable of approximately solving optimization problems with arbitrary objective functions in (B). However, it is important to note that in representing $\beta_{i}$, we introduced two inequality constraints. Here, we introduce the following two approaches:
  • For $s\in[-\frac{1}{2},\frac{1}{2}]$, adding a penalty term
Theorem 4
Example 1 (Sign function)
$\text{sign}(x)$ ($1$ if $x\geq 0$ else $-1$) can be represented as
where $M=\sup|x|$ and $s\in[-\frac{1}{2},\frac{1}{2}]$ is a slack variable.
Appendix C Proof of Convergence of QCGD with Random Quantum Oracle
@@
Appendix D Analysis of Total Error Due to Truncation
@@
$\Delta\text{Hamiltonian}$ $=|E[\sigma|J_{ij},h_{i}]-E[{\sigma^{*}}^{\prime}|J_{ij},h_{i}]|=O(2^{-d}\cdot n^{2}),$ (88)
ΔQUBO $\Delta\text{QUBO Value}$ $=|f(\mathbf{x}|q_{ij})-f({\mathbf{x}^{*}}^{\prime}|q_{ij})|=O(2^{-d}\cdot n^{2}).$ (89)
@@
Appendix E Additional Details of Experimental Results
@@
Appendix F Principles of STE and BinaryConnect
@@
Coherent Ising Machine
As shown in Figure 6, we implement the CIM, an optical system engineered to address combinatorial optimization problems through optical parametric oscillators (OPOs). A continuous wave (CW) laser at 1560 nm initiates the process, feeding into a pump pulse preparation stage. Here, intensity modulators (IM1, IM2, IM3), erbium-doped fiber amplifiers (EDFA1, EDFA2), and filters shape and amplify the light into pump pulses. These pulses enter a second harmonic generation (SHG) unit, producing a frequency-doubled signal, which is then directed into a periodically poled lithium niobate (PPLN) crystal. The PPLN facilitates phase-sensitive amplification, forming a degenerate OPO that generates coherent signal and idler photons. These photons are measured via balanced homodyne detection (BHD), with an FPGA module providing feedback to steer the system toward optimal solutions by mapping the problem onto the Ising model’s energy landscape.
CIM operates by simulating the Ising Hamiltonian, defined as
$H=-\sum_{i,j}J_{ij}\sigma_{i}\sigma_{j}-\sum_{i}h_{i}\sigma_{i},$ (102)
where $\sigma_{i}\in\{1,-1\}$ represents spins, $J_{ij}$ denotes coupling strengths, and $h_{i}$ accounts for external fields. In the PPLN crystal, pump pulses drive parametric amplification, creating optical pulses that encode the spins as two possible phase states. The BHD measures the pulses’ phase and amplitude, determining the system’s energy state. The FPGA uses this data to adjust parameters like pump intensity, effectively minimizing the Hamiltonian through feedback. A chopper and intensity modulators ensure stable pulse circulation in the loop, while the feedback introduces coupling between pulses, mimicking the $J_{ij}\sigma_{i}\sigma_{j}$ interaction terms.
The system’s dynamics allow it to explore the Ising energy landscape efficiently. The optical pulses evolve in parallel, with their interactions governed by the feedback loop, driving the system toward the ground state of the Hamiltonian, which corresponds to the optimization problem’s solution. The amplifiers and filters maintain pulse quality, while the PPLN’s nonlinear efficiency ensures robust signal generation. The BHD’s noise suppression enhances measurement precision, enabling the CIM to handle complex problems by leveraging the coherence and parallelism of optical signals.